# Mantis: A Versatile Vision-Language-Action Model with Disentangled Visual Foresight

### v1

@@ 1 Introduction
Robotic learning has witnessed remarkable progress, particularly in developing robust control strategies for diverse tasks across heterogeneous environments. Among the most promising approaches are Vision-Language-Action (VLA) models [kim2024openvla, brohan2022rt, mu2023embodiedgpt, zawalski2024robotic, zitkovich2023rt, cheang2024gr, team2024octo], which leverage pretrained Vision-Language Models (VLMs) [beyer2024paligemma, bai2025qwen2] to translate linguistic instructions and visual observations into executable robotic actions. Despite these advancements, existing VLA approaches face a fundamental challenge: the low-dimensional action signals can be too sparse to adequately supervise the large VLA model that processes high-dimensional sensory inputs [li2025drivevla]. This mismatch leaves much of the model’s representational capacity underutilized, thereby constraining overall performance.
@@
A remedy is to integrate the visual foresight prediction problem into VLA training, with the model asked to predict dense future visual states in addition to forecasting actions [zhao2025cot, wang2025unified]. However, high-dimensional visual states usually contain redundancies that distract the model from action prediction, resulting in high training costs and slow convergence during downstream fine-tuning [gao2025vla]. Alternative approaches compress visual states into more compact supervisory signals [wen2023any, ye2024latent]. Yet, compression inevitably diminishes subtle variations among visual states that convey fine-grained motions, thereby creating an information bottleneck. Furthermore, existing methods often overlook language supervision, rendering the context understanding and reasoning capabilities of the model unguarded [zhou2025chatvla].
@@
We propose Mantis, a novel VLA model featuring a Disentangled Visual Foresight (DVF) to address these issues. As illustrated in Figure 2, Mantis combines meta queries [pan2025transfer] with a Diffusion Transformer (DiT) [peebles2023scalable] head to predict future visual states. Despite its simplicity, this design can effectively disentangle the tight coupling of action learning and foresight prediction, considering its efficacy in text-to-image generation tasks [pan2025transfer, chen2025blip3]. Furthermore, we advocate feeding the current visual state also to the DiT through a residual connection [he2016deep], which enables the meta queries to automatically capture the inter-frame dynamics that delineate the visual trajectory. Intuitively, these queries extract latent actions which can facilitate more effective explicit action generation; we therefore term them latent-action queries. Then, we use learnable action queries to extract information from both the input and latent-action queries via causal attention. We apply an action head to map the extraction outcomes to explicit actions as in [zhang2025dreamvla].
@@
The disentangled design of Mantis reduces the representation burden on the VLA backbone, thereby freeing up more capacity for language supervision, which helps preserve the model’s semantic understanding and reasoning capabilities after action learning. To minimize the competition between learning signals from various modalities, we develop a progressive training recipe that incrementally incorporates additional modalities for smooth fusion. For inference, Mantis leverages the Temporal Ensemble [zhao2023learning] method for motion stability. Considering its inefficiency issues, we also propose Adaptive Temporal Ensemble (ATE), which dynamically adjusts the ensemble strength according to the motion stability demand of the current timestep. E.g., fine-grained object manipulation can require much higher motion stability than unladen movements. We term the corresponding model variant as Mantis-ATE.
@@
We pretrain Mantis based on three data sources: the SSV2 dataset [goyal2017something] containing 220K human manipulation videos, the DROID dataset [khazatsky2024droid] with 76K robot demonstrations (covering video and action), and image–text pairs drawn from 38 multimodal datasets. We then evaluate Mantis on the LIBERO simulation benchmark [liu2023libero] and in real-world settings. Extensive experiments validate our approach from multiple perspectives: (1) Mantis surpasses several strong baselines [wen2023any, zhao2025cot, cen2025worldvla, bu2025learning, wang2025unified, lv2025f1], achieving 96.7% success on the LIBERO benchmark and showing faster convergence than previous visual foresight approaches [wang2025unified]. (2) Experiments on the Agilex platform confirm Mantis’s robust instruction-following and generalization capabilities. Across three test scenarios, it accurately executes in-domain commands and effectively generalizes to out-of-domain instructions, outperforming the leading open-source VLA model $\pi_{0.5}$ [intelligence2025pi_]. (3) The Mantis-ATE variant reduces inference calls by up to 50% while maintaining comparable task success rates.
@@ 2.1 Vision-Language-Action Models
The rapid evolution of vision-language models (VLMs) [liu2023visual, bai2025qwen2, beyer2024paligemma, team2024chameleon, wu2024vila] has catalyzed the emergence of Vision-Language-Action (VLA) models [zitkovich2023rt, kim2024openvla, brohan2022rt, mu2023embodiedgpt, zawalski2024robotic, cheang2024gr, team2024octo, tian2024predictive, wang2025vla, wen2025tinyvla, li2023vision]. These systems integrate a vision-language backbone with an action prediction component, extending their functional capabilities beyond pure perception. By leveraging the rich perceptual and linguistic representations of pretrained VLMs, VLA models enable robots to interpret and execute human commands with improved adaptability, transcending the limitations of purely reactive control policies. Despite this potential, existing VLA models suffer from limited understanding capabilities. Robot-specific training often overwrites the critical vision-text alignments acquired during pretraining [zhou2025chatvla], degrading instruction-following performance and undermining fundamental reasoning abilities. While some approaches employ explicit language assistance to mitigate this issue [zawalski2024robotic, sun2025emma], they are prone to error accumulation [gao2025vla]. Other methods introduce implicit language supervision through auxiliary losses on language representations [zhou2025chatvla, ji2025robobrain, chen2025training], but lack the visual foresight capabilities necessary for frame-level planning. In contrast, we propose a framework that incorporates visual foresight while simultaneously preserving semantic understanding and reasoning capabilities.
@@ 2.2 Vision-Augmented Action Learning
1) Visual Foresight. This approach enhances action prediction by forecasting future frames, using either explicit or implicit methods. Explicit methods generate future frames through autoregressive discrete image tokens [zhao2025cot, lv2025f1] or learnable query tokens with a vision decoder [wu2023unleashing, zhang2025dreamvla], which then inform action prediction. Implicit methods jointly train the VLA model on video generation and action prediction tasks, establishing an implicit connection between visual forecasting and action prediction [cen2025worldvla, wang2025unified, jiang2025rynnvla]. However, pixel-level forecasting can introduce unnecessary information, distracting the model from action prediction and resulting in high training costs. Moreover, this approach may lead the model to mistakenly associate physical motion with visual appearances, such as texture or lighting changes, potentially resulting in hallucinations [du2023learning].
2) Track Guidance. To mitigate these limitations, several studies compress visual states into compact, control-oriented representations such as keypoint tracks [wen2023any, bharadhwaj2024track2act, bharadhwaj2024gen2act]. These tracks capture essential physical dynamics and facilitate action prediction. Nonetheless, this compression can cause information bottlenecks, and the precision of point-tracking extracted from videos is often limited, leading to higher action prediction errors.
3) Latent Action Supervision. Other approaches employ latent actions to supervise action prediction [ye2024latent, bu2025learning, chen2025moto]. Typically, an action quantization model is first trained to learn discrete latent actions from inter-frame differences. The VLA model is then trained to predict these latent actions before being fine-tuned on robot manipulation data. This strategy leverages the insight that inter-frame dynamics can be represented as action primitives that aid prediction. However, training an additional quantization model increases computational complexity. An overview of these vision-augmentation paradigms is provided in Fig. 1. In contrast to prior work, Mantis decouples visual foresight prediction from the backbone, yielding more compact and accurate auxiliary information for action prediction.
@@ 3.1 Model Overview
This design prevents the VLA backbone from producing redundant visual information directly. Besides, feeding $\mathbf{o}_{t}$ into $\mathcal{D}$ via a residual connection [he2016deep] enables the [LAT] queries to capture inter-frame dynamics that characterize the visual trajectory rather than reconstructing complete frames. These dynamics connect to latent actions—a visual manifestation of explicit robot motions—thereby providing targeted guidance for action prediction.
@@ 3.2 Model Specification
VLM Backbone. Considering the robust understanding and reasoning capabilities of Qwen2.5-VL [bai2025qwen2], we adopt it as our backbone. Qwen2.5-VL natively supports flexible input resolutions, allowing us to assign higher resolution to the primary camera while allocating lower resolution to the wrist camera, which captures less spatial detail.
DVF Head. We employ Sana [xie2024sana] as the DVF head due to its superior performance in text-to-image generation. Sana is a highly efficient DiT that integrates a deep compression autoencoder [chen2024deep]. We define a connector comprising 12 transformer encoder layers and a projection layer that bridges the backbone’s output space with the DiT’s input space, following the Qwen2.5 LLM architecture [yang2025qwen3] while using bidirectional attention. During action inference, the DVF head is omitted to reduce computational overhead, as visual state prediction is not required for robot execution.
Action Head. Following [zhang2025dreamvla], we employ a DiT-based action head for action prediction. Learnable action queries first aggregate information from input and latent-action queries via causal attention, then the action head is used to denoise Gaussian noise into an $n$-step action trajectory.
@@ 3.3 Progressive Training Recipe
Stage 1: Multiple Gap Vision Training. We first train Mantis on videos without action annotations to predict future frames, encouraging the model to infer latent actions from visual dynamics. We use human manipulation videos so that Mantis learns general manipulation skills and broad world knowledge [ye2024latent].
@@ 3.4 Adaptive Temporal Ensemble
During inference, Mantis incorporates the commonly used Temporal Ensemble method [zhao2023learning] to enhance motion stability. Considering its high computational overhead, we also introduce an Adaptive Temporal Ensemble (ATE) strategy that dynamically adjusts the ensemble strength according to the motion stability required at each inference time step.
As shown in Fig. 2 (right), ATE maintains two sets of input vision patches at each time step. (1) Target patches denote image regions that are most relevant to the language instruction. Following [xu2025vla], we compute text-to-vision attention scores from Mantis backbone’s cross-attention module and select the top $\tau_{\text{target}}\%$ of vision tokens with the highest scores. (2) Dynamic patches correspond to regions that exhibit significant visual changes. To detect them, we divide the current and previous input images into patches aligned with the vision tokens, compute the cosine similarity between corresponding patches in pixel space, and select the top $\tau_{\text{dynamic}}\%$ with the lowest similarity values.
@@ 4 Experiments
Spatial Object Goal Long
Avg.
non-vision-
augmented
Diffusion Policy [chi2025diffusion] 78.3 92.5 68.3 50.5 72.4
OpenVLA [kim2024openvla] 84.7 88.4 79.2 53.7 76.5
$\pi_{0}$ [black2024pi_0] 96.8 98.8 95.8 85.2 94.2
$\pi_{0}$-FAST [pertsch2025fast] 96.4 96.8 88.6 60.2 85.5
NORA [hung2025nora] 92.2 95.4 89.4 74.6 87.9
vision-
augmented
ATM [wen2023any] 68.5 68.0 77.8 39.3 63.4
CoT-VLA [zhao2025cot] 87.5 91.6 87.6 69.0 81.1
WorldVLA [cen2025worldvla] 87.6 96.2 83.4 60.0 81.8
UniVLA [bu2025learning] 96.5 96.8 95.6 92.0 95.2
UnifiedVLA [wang2025unified] 95.4 98.8 93.6 94.0 95.5
DreamVLA [zhang2025dreamvla] 97.5 94.0 89.5 89.5 92.6
$\mathcal{F}_{1}$ [lv2025f1] 98.2 97.8 95.4 91.3 95.7
@@ 4.1 Implementation Details
Basic configuration. Mantis comprises 5.8 billion parameters in total: 3.7B in the backbone, 1.4B in the DVF head, 0.3B in the action head, and 0.3B in the VAE. We set the number of [LAT] to 9, [ACT] to 6, and [GAP] to $6\times 3$ corresponding to time step gaps from 1 to 6. The diffusion process uses 30 steps for the DVF head and 10 steps for the action head. Training adopts AdamW [loshchilov2017decoupled] with 0.1 weight decay and 0.5 gradient clipping, and leverages DeepSpeed [rasley2020deepspeed] for efficient distributed training.
Pretraining Setup. In Stage 1, the model is pretrained on the SSV2 dataset [goyal2017something], which contains approximately 220K videos of human manipulations. The time step gap is randomly sampled between 1 and 6. In Stage 2, the model is pretrained on the DROID dataset [khazatsky2024droid], consisting of 76K robot episodes. The vision loss weight $\alpha$ is set to 0.1. In Stage 3, we introduce language supervision by jointly training on 38 multimodal datasets and DROID for 1.5 epochs.
Finetuning Setup. For downstream finetuning on the LIBERO benchmark [liu2023libero], we use the same learning rate configuration as in pretraining stages 1 and 2 and train for 30 epochs without language supervision. The vision loss weight is set to $\alpha=0.1$. The checkpoint with the highest validation success rate (SR) is selected for final evaluation. Further implementation details are provided in Appendix B.
@@ 4.2 Simulation Experiments
Figure 5: Convergence speed comparison. Compared with traditional visual foresight methods such as UnifiedVLA [wang2025unified], Mantis achieves significantly faster convergence speed, underscoring the necessity of decoupling foresight prediction from action learning.
@@
We evaluate Mantis on the widely adopted LIBERO benchmark [liu2023libero], which comprises four task suites: Spatial, Object, Goal, and Long. Each suite contains 10 tasks, with 50 trials per task to reduce random variation. Performance is measured using Success Rate (SR) from 0 to 100, where higher is better. For comprehensive comparison, we select high-performance VLA models both with [wen2023any, zhao2025cot, cen2025worldvla, bu2025learning, wang2025unified, zhang2025dreamvla, lv2025f1] and without [chi2025diffusion, kim2024openvla, black2024pi_0, pertsch2025fast, hung2025nora] vision augmentation as baselines. We also compare the convergence speed of Mantis against four vision-augmented baselines.
Main Results. As shown in Table 1, Mantis exhibits superior performance on 3 of 4 task suites and attains the highest average SR, outperforming previous vision-augmented and non-vision-augmented methods. This underscores the efficacy of leveraging DVF to enhance action prediction. Furthermore, vision-augmented methods mostly outperform their non-vision-augmented counterparts, corroborating the findings in [li2025drivevla] that dense visual states effectively complement sparse action signals. Notably, ATM [wen2023any] exhibits inferior performance, which we attribute to the limited accuracy of point trajectories extracted via video tracking methods, resulting in accumulated errors.
Convergence Speed. We evaluate the convergence speed of Mantis and four baseline methods representing different learning paradigms: non-vision-augmentation (OpenVLA [kim2024openvla]), visual foresight (UnifiedVLA [wang2025unified]), track guidance (ATM [wen2023any]), and latent action supervision (UniVLA [bu2025learning]). Each model is fine-tuned from its pretrained checkpoint for 20 epochs on the LIBERO Spatial suite following the prescribed configuration. Evaluation is conducted at each epoch, and the results are presented as an epoch-SR line graph in Fig. 5. As illustrated, Mantis demonstrates a relatively fast convergence speed, comparable to the non-vision-augmented OpenVLA and the latent action supervision method UniVLA. In contrast, the entangled visual foresight approach, UnifiedVLA, converges the slowest, maintaining a success rate of zero during the first ten epochs. This observation highlights the necessity of decoupling foresight prediction from action learning in order to achieve efficient optimization.

### v2

@@ 1 Introduction
Robotic learning has witnessed remarkable progress, particularly in developing robust control strategies for diverse tasks across heterogeneous environments. Among the most promising approaches are Vision-Language-Action (VLA) models [25, 7, 33, 53, 58, 10, 40], which leverage pretrained Vision-Language Models (VLMs) [3, 2] to translate linguistic instructions and visual observations into executable robotic actions. Despite these advancements, existing VLA approaches face a fundamental challenge: the low-dimensional action signals can be too sparse to adequately supervise the large VLA model that processes high-dimensional sensory inputs [27]. This mismatch leaves much of the model’s representational capacity underutilized, thereby constraining overall performance.
@@
A remedy is to integrate the visual foresight prediction problem into VLA training, with the model asked to predict dense future visual states in addition to forecasting actions [55, 43]. However, high-dimensional visual states usually contain redundancies that distract the model from action prediction, resulting in high training costs and slow convergence during downstream fine-tuning [17]. Alternative approaches compress visual states into more compact supervisory signals [44, 51]. Yet, compression inevitably diminishes subtle variations among visual states that convey fine-grained motions, thereby creating an information bottleneck. Furthermore, existing methods often overlook language supervision, rendering the context understanding and reasoning capabilities of the model unguarded [57].
@@
We propose Mantis, a novel VLA model featuring a Disentangled Visual Foresight (DVF) to address these issues. As illustrated in Figure 2, Mantis combines meta queries [34] with a Diffusion Transformer (DiT) [35] head to predict future visual states. Despite its simplicity, this design can effectively disentangle the tight coupling of action learning and foresight prediction, considering its efficacy in text-to-image generation tasks [34, 12]. Furthermore, we advocate feeding the current visual state also to the DiT through a residual connection [19], which enables the meta queries to automatically capture the inter-frame dynamics that delineate the visual trajectory. Intuitively, these queries extract latent actions which can facilitate more effective explicit action generation; we therefore term them latent-action queries. Then, we use learnable action queries to extract information from both the input and latent-action queries via causal attention. We apply an action head to map the extraction outcomes to explicit actions as in [54].
@@
The disentangled design of Mantis reduces the representation burden on the VLA backbone, thereby freeing up more capacity for language supervision, which helps preserve the model’s semantic understanding and reasoning capabilities after action learning. To minimize the competition between learning signals from various modalities, we develop a progressive training recipe that incrementally incorporates additional modalities for smooth fusion. For inference, Mantis leverages the Temporal Ensemble [56] method for motion stability. Considering its inefficiency issues, we also propose Adaptive Temporal Ensemble (ATE), which dynamically adjusts the ensemble strength according to the motion stability demand of the current timestep. E.g., fine-grained object manipulation can require much higher motion stability than unladen movements. We term the corresponding model variant as Mantis-ATE.
@@
We pretrain Mantis based on three data sources: the SSV2 dataset [18] containing 220K human manipulation videos, the DROID dataset [24] with 76K robot demonstrations (covering video and action), and image–text pairs drawn from 38 multimodal datasets. We then evaluate Mantis on the LIBERO simulation benchmark [28] and in real-world settings. Extensive experiments validate our approach from multiple perspectives: (1) Mantis surpasses several strong baselines [44, 55, 9, 8, 43, 32], achieving 96.7% success on the LIBERO benchmark and showing faster convergence than previous visual foresight approaches [43]. (2) Experiments on the Agilex platform confirm Mantis’s robust instruction-following and generalization capabilities. Across three test scenarios, it accurately executes in-domain commands and effectively generalizes to out-of-domain instructions, outperforming the leading open-source VLA model $\pi_{0.5}$ [21]. (3) The Mantis-ATE variant reduces inference calls by up to 50% while maintaining comparable task success rates.
@@ 2.1 Vision-Language-Action Models
The rapid evolution of vision-language models (VLMs) [29, 2, 3, 39, 47] has catalyzed the emergence of Vision-Language-Action (VLA) models [58, 25, 7, 33, 53, 10, 40, 41, 42, 45, 26]. These systems integrate a vision-language backbone with an action prediction component, extending their functional capabilities beyond pure perception. By leveraging the rich perceptual and linguistic representations of pretrained VLMs, VLA models enable robots to interpret and execute human commands with improved adaptability, transcending the limitations of purely reactive control policies. Despite this potential, existing VLA models suffer from limited understanding capabilities. Robot-specific training often overwrites the critical vision-text alignments acquired during pretraining [57], degrading instruction-following performance and undermining fundamental reasoning abilities. While some approaches employ explicit language assistance to mitigate this issue [53, 38], they are prone to error accumulation [17]. Other methods introduce implicit language supervision through auxiliary losses on language representations [57, 22, 13], but lack the visual foresight capabilities necessary for frame-level planning. In contrast, we propose a framework that incorporates visual foresight while simultaneously preserving semantic understanding and reasoning capabilities.
@@ 2.2 Vision-Augmented Action Learning
1) Visual Foresight. This approach enhances action prediction by forecasting future frames, using either explicit or implicit methods. Explicit methods generate future frames through autoregressive discrete image tokens [55, 32] or learnable query tokens with a vision decoder [46, 54], which then inform action prediction. Implicit methods jointly train the VLA model on video generation and action prediction tasks, establishing an implicit connection between visual forecasting and action prediction [9, 43, 23]. However, pixel-level forecasting can introduce unnecessary information, distracting the model from action prediction and resulting in high training costs. Moreover, this approach may lead the model to mistakenly associate physical motion with visual appearances, such as texture or lighting changes, potentially resulting in hallucinations [16].
2) Track Guidance. To mitigate these limitations, several studies compress visual states into compact, control-oriented representations such as keypoint tracks [44, 5, 4]. These tracks capture essential physical dynamics and facilitate action prediction. Nonetheless, this compression can cause information bottlenecks, and the precision of point-tracking extracted from videos is often limited, leading to higher action prediction errors.
3) Latent Action Supervision. Other approaches employ latent actions to supervise action prediction [51, 8, 14]. Typically, an action quantization model is first trained to learn discrete latent actions from inter-frame differences. The VLA model is then trained to predict these latent actions before being fine-tuned on robot manipulation data. This strategy leverages the insight that inter-frame dynamics can be represented as action primitives that aid prediction. However, training an additional quantization model increases computational complexity. An overview of these vision-augmentation paradigms is provided in Fig. 1. In contrast to prior work, Mantis decouples visual foresight prediction from the backbone, yielding more compact and accurate auxiliary information for action prediction.
@@ 3.1 Model Overview
This design prevents the VLA backbone from producing redundant visual information directly. Besides, feeding $\mathbf{o}_{t}$ into $\mathcal{D}$ via a residual connection [19] enables the [LAT] queries to capture inter-frame dynamics that characterize the visual trajectory rather than reconstructing complete frames. These dynamics connect to latent actions—a visual manifestation of explicit robot motions—thereby providing targeted guidance for action prediction.
@@ 3.2 Model Specification
VLM Backbone. Considering the robust understanding and reasoning capabilities of Qwen2.5-VL [2], we adopt it as our backbone. Qwen2.5-VL natively supports flexible input resolutions, allowing us to assign higher resolution to the primary camera while allocating lower resolution to the wrist camera, which captures less spatial detail.
DVF Head. We employ Sana [48] as the DVF head due to its superior performance in text-to-image generation. Sana is a highly efficient DiT that integrates a deep compression autoencoder [11]. We define a connector comprising 12 transformer encoder layers and a projection layer that bridges the backbone’s output space with the DiT’s input space, following the Qwen2.5 LLM architecture [50] while using bidirectional attention. During action inference, the DVF head is omitted to reduce computational overhead, as visual state prediction is not required for robot execution.
Action Head. Following [54], we employ a DiT-based action head for action prediction. Learnable action queries first aggregate information from input and latent-action queries via causal attention, then the action head is used to denoise Gaussian noise into an $n$-step action trajectory.
@@ 3.3 Progressive Training Recipe
Stage 1: Multiple Gap Vision Training. We first train Mantis on videos without action annotations to predict future frames, encouraging the model to infer latent actions from visual dynamics. We use human manipulation videos so that Mantis learns general manipulation skills and broad world knowledge [51].
@@ 3.4 Adaptive Temporal Ensemble
During inference, Mantis incorporates the commonly used Temporal Ensemble method [56] to enhance motion stability. Considering its high computational overhead, we also introduce an Adaptive Temporal Ensemble (ATE) strategy that dynamically adjusts the ensemble strength according to the motion stability required at each inference time step.
As shown in Fig. 2 (right), ATE maintains two sets of input vision patches at each time step. (1) Target patches denote image regions that are most relevant to the language instruction. Following [49], we compute text-to-vision attention scores from Mantis backbone’s cross-attention module and select the top $\tau_{\text{target}}\%$ of vision tokens with the highest scores. (2) Dynamic patches correspond to regions that exhibit significant visual changes. To detect them, we divide the current and previous input images into patches aligned with the vision tokens, compute the cosine similarity between corresponding patches in pixel space, and select the top $\tau_{\text{dynamic}}\%$ with the lowest similarity values.
@@ 4 Experiments
Spatial Object Goal Long Avg.
non-vision- augmented Diffusion Policy [15] 78.3 92.5 68.3 50.5 72.4
OpenVLA [25] 84.7 88.4 79.2 53.7 76.5
$\pi_{0}$ [6] 96.8 98.8 95.8 85.2 94.2
$\pi_{0}$-FAST [36] 96.4 96.8 88.6 60.2 85.5
NORA [20] 92.2 95.4 89.4 74.6 87.9
vision- augmented ATM [44] 68.5 68.0 77.8 39.3 63.4
CoT-VLA [55] 87.5 91.6 87.6 69.0 81.1
WorldVLA [9] 87.6 96.2 83.4 60.0 81.8
UniVLA [8] 96.5 96.8 95.6 92.0 95.2
UnifiedVLA [43] 95.4 98.8 93.6 94.0 95.5
DreamVLA [54] 97.5 94.0 89.5 89.5 92.6
$\mathcal{F}_{1}$ [32] 98.2 97.8 95.4 91.3 95.7
@@ 4.1 Implementation Details
Basic configuration. Mantis comprises 5.8 billion parameters in total: 3.7B in the backbone, 1.4B in the DVF head, 0.3B in the action head, and 0.3B in the VAE. We set the number of [LAT] to 9, [ACT] to 6, and [GAP] to $6\times 3$ corresponding to time step gaps from 1 to 6. The diffusion process uses 30 steps for the DVF head and 10 steps for the action head. Training adopts AdamW [31] with 0.1 weight decay and 0.5 gradient clipping, and leverages DeepSpeed [37] for efficient distributed training.
Pretraining Setup. In Stage 1, the model is pretrained on the SSV2 dataset [18], which contains approximately 220K videos of human manipulations. The time step gap is randomly sampled between 1 and 6. In Stage 2, the model is pretrained on the DROID dataset [24], consisting of 76K robot episodes. The vision loss weight $\alpha$ is set to 0.1. In Stage 3, we introduce language supervision by jointly training on 38 multimodal datasets and DROID for 1.5 epochs.
Finetuning Setup. For downstream finetuning on the LIBERO benchmark [28], we use the same learning rate configuration as in pretraining stages 1 and 2 and train for 30 epochs without language supervision. The vision loss weight is set to $\alpha=0.1$. The checkpoint with the highest validation success rate (SR) is selected for final evaluation. Further implementation details are provided in Appendix B.
@@ 4.2 Simulation Experiments
Figure 5: Convergence speed comparison. Compared with traditional visual foresight methods such as UnifiedVLA [43], Mantis achieves significantly faster convergence speed, underscoring the necessity of decoupling foresight prediction from action learning.
@@
We evaluate Mantis on the widely adopted LIBERO benchmark [28], which comprises four task suites: Spatial, Object, Goal, and Long. Each suite contains 10 tasks, with 50 trials per task to reduce random variation. Performance is measured using Success Rate (SR) from 0 to 100, where higher is better. For comprehensive comparison, we select high-performance VLA models both with [44, 55, 9, 8, 43, 54, 32] and without [15, 25, 6, 36, 20] vision augmentation as baselines. We also compare the convergence speed of Mantis against four vision-augmented baselines.
Main Results. As shown in Table 1, Mantis exhibits superior performance on 3 of 4 task suites and attains the highest average SR, outperforming previous vision-augmented and non-vision-augmented methods. This underscores the efficacy of leveraging DVF to enhance action prediction. Furthermore, vision-augmented methods mostly outperform their non-vision-augmented counterparts, corroborating the findings in [27] that dense visual states effectively complement sparse action signals. Notably, ATM [44] exhibits inferior performance, which we attribute to the limited accuracy of point trajectories extracted via video tracking methods, resulting in accumulated errors.
Convergence Speed. We evaluate the convergence speed of Mantis and four baseline methods representing different learning paradigms: non-vision-augmentation (OpenVLA [25]), visual foresight (UnifiedVLA [43]), track guidance (ATM [44]), and latent action supervision (UniVLA [8]). Each model is fine-tuned from its pretrained checkpoint for 20 epochs on the LIBERO Spatial suite following the prescribed configuration. Evaluation is conducted at each epoch, and the results are presented as an epoch-SR line graph in Fig. 5. As illustrated, Mantis demonstrates a relatively fast convergence speed, comparable to the non-vision-augmented OpenVLA and the latent action supervision method UniVLA. In contrast, the entangled visual foresight approach, UnifiedVLA, converges the slowest, maintaining a success rate of zero during the first ten epochs. This observation highlights the necessity of decoupling foresight prediction from action learning in order to achieve efficient optimization.